# Detection of Non-recorded Word Senses in English and Swedish

### v1

@@ Abstract
This study addresses the task of Unknown Sense Detection in English and Swedish. The primary objective of this task is to determine whether the meaning of a particular word usage is documented in a dictionary or not. For this purpose, sense entries are compared with word usages from modern and historical corpora using a pre-trained Word-in-Context embedder that allows us to model this task in a few-shot scenario. Additionally, we use human annotations to adapt and evaluate our models. Compared to a random sample from a corpus, our model is able to considerably increase the detected number of word usages with non-recorded senses.
@@
Simon Hengchen iguanodon.ai & University of Geneva simon@iguanodon.ai                        Dominik Schlechtweg IMS, University of Stuttgart schlecdk@ims.uni-stuttgart.de
@@ 1 Introduction
In this paper, we investigate systems that discover missing dictionary sense entries in modern English and Swedish dictionaries by comparing target word usages from reference corpora to the dictionary entries for the target word. The basic task to solve is to decide whether the sense of a given word usage is covered by any sense in the dictionary entry of the target word or not. The major difficulty is that a dictionary entry, despite containing a good amount of lexical information, does not provide enough context in order to train a standard Word Sense Disambiguation (WSD) system. Depending on the dictionary, the senses of an entry are sometimes only covered by the definition or a few example sentences. We are addressing this problem by using a pre-trained Word-in-Context (WiC) model to create contextualized word embeddings (Cassotti et al., 2023) and further comparing different ways of maximizing the quality of information we utilize from the dictionary entries. Although we are not the first to address this problem (Erk, 2006; Lau et al., 2012), none of the related approaches utilizes WiC models to solve the task in a realistic few-shot scenario. Moreover, our model can even be applied to multilingual applications. Furthermore, we use a human annotation approach for a high-quality evaluation of our model’s performance.
@@ 2.1 Word Sense Disambiguation
Word Sense Disambiguation (WSD) is a classical, yet unsolved task in NLP and has been studied for decades (Weaver, 1949/1955). Navigli (2009) describes the task of WSD as “computationally identifying the meaning of a word by its use in a particular context” (p. 1). A word sequence $s$ that contains a particular target word $w$, at the position $i$ we call a word usage of $w$ and represent it by the tuple $u=(s,i)$. Navigli (2009, p. 4) defines a word sense as a commonly accepted meaning of a word. Which meaning is expressed in a particular word usage depends on the context of use. Let $\Phi=\{\phi_{1},\phi_{2},\dots,\phi_{k}\}$ be a set of predefined senses and $f:U\rightarrow\Phi$ be a mapping from usages to senses given by a lexical resource. The task is to find mapping $f$. So by definition, WSD can be considered a multiclass classification task.
@@
In order to correctly classify a word usage based on a predefined set of senses, the sense has to be part of the set in the first place. Building such sense inventory with good coverage is strongly dependent on the available data. Traditional WSD models use large amounts of manually annotated data (multi-shot) to train one classifier per word to disambiguate its word usages (Burchardt et al., 2009). Since manually labeled data sets are very time- and resource-consuming to produce, they are usually of a limited size and do not cover the full lexicon of the language. While these classifier models generally deliver good results on a restricted set of words, they quickly suffer from the problem of data sparseness (Navigli, 2009, p. 4, p. 16). More realistic WSD models therefore try to reduce the dependency on manually labeled training data. Building such models makes it possible to extract the little training data needed (few-shot) from limited lexical resources like a dictionary which maps headwords [^1] In the remainder of this paper, we use ‘headword’ when referring to an entry in a lexical resource, and ‘lemma’ as the canonical form of a given word. The two concepts greatly overlap. to sets of sense entries. A sense entry is usually structured, providing a sense definition (gloss), a number of example usages of the headword in that sense and some meta information such as the etymology and part-of-speech (see Appendix B for an example). Early few-shot computational approaches to WSD make use of such dictionaries and try to disambiguate word usages by simply counting overlapping words between eligible sense definitions of the target word and sense definitions of nearby words (Lesk, 1986). Newer models often use pre-trained contextualized embedders (Devlin et al., 2019; Peters et al., 2018a) to calculate vector representations for target usages and senses (Section 2.4), which can then be compared, e.g. with a similarity metric (Scarlini et al., 2020; Hu et al., 2019; Kumar et al., 2019). Rachinskiy and Arefyev (2022) develop a model training only one classifier for all words in their data. In this way, knowledge between words can be shared, e.g., from frequent to infrequent ones (Kågebäck and Salomonsson, 2016; Chen et al., 2021).
@@ 2.4 Contextualized Word Meaning Representations
BERT is one such contextualized embedding model that generates a high-dimensional vector representation for a given context. The model architecture is based on the original multi-layer bidirectional Transformer encoder by Vaswani et al. (2017). BERT-embeddings find use in various NLP tasks (Devlin et al., 2019). SentenceBERT (SBERT) extends BERT, using Siamese and triplet networks to produce meaningful sense embeddings for a single sentence (Reimers and Gurevych, 2019). This approach is highly efficient in the sense that it dramatically reduces inference time while introducing no loss of performance. Building on SBERT, the recent XL-LEXEME (Cassotti et al., 2023) is based on the pre-trained SBERT and fine-tuned on human-labeled Word-in-Context (WiC, Pilehvar and Camacho-Collados, 2019) data. It has been shown to perform extremely well on lexical semantics tasks such as lexical semantic change detection (Schlechtweg et al., 2020) by highlighting target words in sentences.
@@ 3.1 Corpora
Since our goal is to support dictionary maintenance, we have to make sure to create a realistic scenario. Use of language changes over time, which necessarily leads to the emergence of sense entry gaps in dictionaries. Unrecorded word senses can occur for two main reasons: either they are old senses that were left unrecorded when the dictionary was created, or they are novel senses that emerged due to language change. The former are likely to be found in historical data while the novel senses more likely occur in modern corpora. To evaluate our models on both cases, we decided to use modern and historical corpora. Table 2 shows an overview of our corpora. A precise statistical description of the modern corpora can be found on the website of Wortschatz Leipzig. [^2] https://cls.corpora.uni-leipzig.de/
@@ 3.2 Lexical resources
For Swedish, we use a processed data dump of Svensk ordbok (SO), [^3] https://svenska.se/so/. Please refer to the Limitations section. the main contemporary dictionary for the Swedish language, created and maintained by the University of Gothenburg (Allén, 1981). It is structured in the classic style of a dictionary in that it lists senses under a headword. In order to use a standardized form for such a headword, Allén (1981) applies the lemma-lexeme model and groups different inflections and forms of words by their lemma. In this way, our processed data dump of SO stores a total of 68,000 senses for over 41,500 headwords. The majority of the senses are described by a gloss, in some cases extended by a secondary gloss. [^4] We only made use of the secondary gloss if no standard gloss was given. The secondary gloss gives context to the primary gloss, if existing. All senses are followed by example usages illustrating how the headword is used.
The two dictionaries [^5] We use the term “dictionary” for WordNet here and in the remainder of this paper after having transformed it into a dictionary-like structure (see above). represent the senses of words, describe them using both gloss and example usages, and group them by their headword. Both dictionaries differ in that their entries have gaps in different parts of their sense entries. While WordNet senses are fully covered by gloss but only have partial example usages, it is the other way around in the Swedish dictionary. A direct comparison of the statistics of both dictionaries is shown in Table 3.
@@ 4 Models
All our models have the same basic structure but differ in the type of data used and the embedding comparison methodology. Essentially, they first use the information given in the dictionary to create a vector representation for each existing sense. The type of information used (e.g., gloss, example usage) and how it is processed (e.g., inserting headwords if missing) are treated as model hyperparameters, as listed below. The models then create a usage embedding for the word usage to be examined and calculate the similarity to all sense embeddings of the associated headword. If the calculated similarity score between the usage embedding and the closest sense embedding is below a specific threshold, the model predicts that the word usage is the occurrence of an unknown sense and therefore marks it as unassigned. [^6] Find our code at: https://github.com/ChangeIsKey/non-recorded-sense-detection. We decided to use XL-LEXEME (see Section 2) to vectorize both target usages and senses because of its good performance in Lexical Semantic Change Detection. This task involves the detection of word sense changes and thus XL-LEXEME can be expected to encode word sense information.
Each model’s identifier is made up of a different choice of the hyperparameters [^7] The total computational cost of this hyperparameter search is roughly 100 hours of GPU time., which will be explained in more detail in the following paragraphs:
@@ 5 Annotation
The annotators are presented instances consisting of a word usage where the target word is marked and a sense gloss that may or may not describe this usage. Additionally, they are shown the list of all possible sense glosses for the target word. The annotation approach takes inspiration from Erk et al. (2013)’s WSsim method by conducting an individual assessment of all senses for a usage, but differs in that we ask annotators for a binary classification into sense gloss fits (label “1”) and sense gloss does not fit (label “0”), instead of a five-level rating. The annotators also always have the option to choose that no specification is possible (label “-”). They are encouraged to leave a comment if this is the case. An example of an annotation instance for a usage of the word relative can be found in Appendix A.
@@ 5.1 Phase I: Random sample
In each corpus, we search word usages of a random sample of 3,000 headwords until we find at least one usage for 150 of them. We keep at most five usages for each headword, chosen at random. Lastly, we combine the word usages from the modern and the historical samples. This results in a set of word usages that is approximately equally distributed between modern and historical corpora, which can be linked to a headword in the dictionary. Combining each word usage with each eligible sense returns approximately 1,200 annotation instances, distributed over 500–700 usages, as listed in Table 5. These are assessed by the three annotators. For WordNet, only primary synsets are considered in the first annotation phase.
@@
We report inter-annotator agreement scores measured by Krippendorf’s alpha (Krippendorff, 2018) in Table 8 in Appendix F. Agreement is low to moderate: In English it ranges from 0.273 to 0.561, while Swedish has more agreement (minimum 0.317, maximum 0.655). The pairwise comparison reveals that annotator A1 of English stands out in particular and the observation holds for both modern and historical usages. [^11] This annotator was excluded from the second annotation phase. For Swedish, the annotators are more in agreement, especially on the modern data.
@@ 5.2 Phase II: Model prediction
Find a summary of the annotation results in Table 5. In comparison to the first phase, the second phase yields considerably higher shares of unassigned usages for both languages, which will be discussed in more detail in Section 6.2.
For English, the annotator-agreement on the full data is 0.384, as displayed in Table 9 in Appendix F. The agreement is comparable to the first phase. However, note that we increased the number of senses the annotators can choose from by including the number of non-primary synsets. There is considerably higher consensus among the Swedish annotators (0.56). For English, agreement on historical instances is slightly higher than on the modern instances. This contrasts with phase I. For Swedish, it is considerably worse on historical instances, as in phase I.
@@ 6.1 Phase I: Model selection
We select hyper-parameters for our models on the randomly sampled and annotated data from Phase I (see Section 5.1). In order not to overfit them on the limited data, we assess them using k-fold cross-validation. We perform a total of 10 rounds of 5-fold cross-validation for every model. Each round consists of different manipulations of the gold sense assignments. In each round, we randomly mask a set of assigned senses as unassigned in order to simulate unknown senses (cf. the method described in Erk, 2006) because the number the number of naturally unassigned senses was very low. [^12] Unassigned senses are left untouched and are included in the evaluation data. The models have to know at least one sense of a headword to examine its usages (see Section 4). Therefore, under certain conditions, masking can lead to loss of data as it could for example mask the only complete (i.e., providing sufficient information) sense of a headword. In order to prevent this, we have to consider three special cases:
  1. 1.
    Senses of headwords that have no complete sense are excluded from the model evaluation since its word usages cannot be compared against any sense vector.
  2. 2.
    Incomplete senses are always masked as a model could never assign this sense.
  3. 3.
    Senses that are the only complete sense of the headword are always unmasked as masking them would exclude it from evaluation.
Note that headwords of all remaining senses have multiple complete senses and can therefore be randomly masked.
We use the same simulated data for all models of a language, even though the gloss models in English and the example models in Swedish do not suffer from limited dictionary completeness. To further ensure the models’ capability to examine every word usage after masking, we randomly select one sense for each headword to exclude before masking any remaining. This prevents us from accidentally masking all senses of a headword.
@@
Applying the randomization ten times leaves us with ten different data sets of 5-fold subdivided usages. In each round for each fold $k$, the model uses the data of the remaining four training-folds to determine a threshold $\in\{0.0,0.01,0.02,\dots,1.0\}$ (see Section 4) for the respective similarity measure, so that the $F_{0.3}$-score is maximized, giving more importance to precision over recall. [^13] Find a discussion of the evaluation metric selection in Appendix C. Using this threshold, the model then predicts on the test-fold $k$.

### v2

@@ Abstract
This study addresses the task of Unknown Sense Detection in English and Swedish. The primary objective of this task is to determine whether the meaning of a particular word usage is documented in a dictionary or not. For this purpose, sense entries are compared with word usages from modern and historical corpora using a pre-trained Word-in-Context embedder that allows us to model this task in a few-shot scenario. Additionally, we use human annotations on the target corpora to adapt hyperparameters and evaluate our models using 5-fold cross-validation. Compared to a random sample from a corpus, our model is able to considerably increase the detected number of word usages with non-recorded senses.
@@
Simon Hengchen iguanodon.ai & University de Genève simon@iguanodon.ai                        Dominik Schlechtweg IMS, University of Stuttgart schlecdk@ims.uni-stuttgart.de
@@ 1 Introduction
In this paper, we investigate systems that discover missing dictionary sense entries in modern English and Swedish dictionaries by comparing target word usages from reference corpora to the dictionary entries for the target word. The basic task to solve is to decide whether the sense of a given word usage is covered by any sense in the dictionary entry of the target word or not. The major difficulty is that a dictionary entry, despite containing a good amount of lexical information, does not provide enough context in order to train a standard Word Sense Disambiguation (WSD) system. Depending on the dictionary, the senses of an entry are sometimes only covered by the definition or a few example sentences. We are addressing this problem by using a pre-trained Word-in-Context (WiC) model to create contextualized word embeddings (Cassotti et al., 2023) and further comparing different ways of maximizing the quality of information we utilize from the dictionary entries. Although we are not the first to address this problem (Erk, 2006; Lau et al., 2012), none of the related approaches utilizes WiC models to solve the task in a realistic scenario: Erk does not operate in a few-shot scenario while Lau et al. do not automatically compare induced senses to a dictionary (but only between corpora). Moreover, our model can even be applied to multilingual applications. Furthermore, we use a human annotation approach for a high-quality evaluation of our model’s performance.
@@ 2.1 Word Sense Disambiguation
Word Sense Disambiguation (WSD) is a classical, yet unsolved task in NLP and has been studied for decades (Weaver, 1949/1955; Resnik and Yarowsky, 1997). Navigli (2009) describes the task of WSD as “computationally identifying the meaning of a word by its use in a particular context” (p. 1). A word sequence $s$ that contains a particular target word $w$, at the position $i$ we call a word usage of $w$ and represent it by the tuple $u=(s,i)$. Navigli (2009, p. 4) defines a word sense as a commonly accepted meaning of a word. Which meaning is expressed in a particular word usage depends on the context of use. Let $\Phi=\{\phi_{1},\phi_{2},\dots,\phi_{k}\}$ be a set of predefined senses and $f:U\rightarrow\Phi$ be a mapping from usages to senses given by a lexical resource. The task is to find mapping $f$. So by definition, WSD can be considered a multiclass classification task.
@@
In order to correctly classify a word usage based on a predefined set of senses, the sense has to be part of the set in the first place. Building such sense inventory with good coverage is strongly dependent on the available data. Traditional WSD models use large amounts of manually annotated data (multi-shot) to train one classifier per word to disambiguate its word usages (Burchardt et al., 2009). Since manually labeled data sets are very time- and resource-consuming to produce, they are usually of a limited size and do not cover the full lexicon of the language. While these classifier models generally deliver good results on a restricted set of words, they quickly suffer from the problem of data sparseness (Navigli, 2009, p. 4, p. 16). More realistic WSD models therefore try to reduce the dependency on manually labeled training data. Building such models makes it possible to extract the little training data needed (few-shot) from limited lexical resources like a dictionary which maps headwords [^1] In the remainder of this paper, we use ‘headword’ when referring to an entry in a lexical resource, and ‘lemma’ as the canonical form of a given word. The two concepts greatly overlap. to sets of sense entries. A sense entry is usually structured, providing a sense definition (gloss), a number of example usages of the headword in that sense and some meta information such as the etymology and part-of-speech (see Appendix A for an example). Early few-shot computational approaches to WSD make use of such dictionaries and try to disambiguate word usages by simply counting overlapping words between eligible sense definitions of the target word and sense definitions of nearby words (Lesk, 1986). Newer models often use pre-trained contextualized embedders (Devlin et al., 2019; Peters et al., 2018a) to calculate vector representations for target usages and senses (Section 2.4), which can then be compared, e.g. with a similarity metric (Scarlini et al., 2020; Hu et al., 2019; Kumar et al., 2019). Rachinskiy and Arefyev (2022) develop a model training only one classifier for all words in their data. In this way, knowledge between words can be shared, e.g., from frequent to infrequent ones (Kågebäck and Salomonsson, 2016; Chen et al., 2021).
@@ 2.4 Contextualized Word Meaning Representations
BERT is one such contextualized embedding model that generates a high-dimensional vector representation for a given context. The model architecture is based on the original multi-layer bidirectional Transformer cross-encoder by Vaswani et al. (2017). BERT-embeddings find use in various NLP tasks (Devlin et al., 2019). SentenceBERT (SBERT) is a modification of BERT overcoming limitations of cross-encoders using a bi-encoder in a Siamese Network, i.e., input sentences are encoded separately while the weights of the underlying networks are shared (Reimers and Gurevych, 2019). The recent XL-LEXEME model (Cassotti et al., 2023) adapts the SBERT architecture to the Word-in-Context task (WiC, Pilehvar and Camacho-Collados, 2019) by giving prominence to the target word enclosing it into special tokens. The model was trained to minimize the contrastive loss with cosine distance on the multilingual WiC dataset (Pilehvar and Camacho-Collados, 2019), using XLMR (Conneau et al., 2019) as the underlying language model. It has been shown to perform extremely well on lexical semantic tasks such as lexical semantic change detection (Schlechtweg et al., 2020).
@@ 3.1 Corpora
Since our goal is to support dictionary maintenance, we have to make sure to create a realistic scenario. Use of language changes over time, which necessarily leads to the emergence of sense entry gaps or outdated sense entries in dictionaries. Unrecorded word senses can occur for two main reasons: either they are old senses that were left unrecorded when the dictionary was created, or they are novel senses that emerged due to language change. The former are likely to be found in historical data while the novel senses more likely occur in modern corpora. To evaluate our models on both cases, we decided to use modern and historical corpora. Table 2 shows an overview of our corpora. A precise statistical description of the modern corpora can be found on the website of Wortschatz Leipzig. [^2] https://cls.corpora.uni-leipzig.de/
@@ 3.2 Lexical resources
For Swedish, we use a processed data dump of Svensk ordbok (SO), [^3] https://svenska.se/so/. Please refer to the Limitations section. the main contemporary dictionary for the Swedish language, created and maintained by the University of Gothenburg (Allén, 1981). It is structured in the classic style of a dictionary in that it lists senses under a headword. In order to use a standardized form for such a headword, Allén (1981) applies the lemma-lexeme model and groups different inflections and forms of words by their lemma. In this way, our processed data dump of SO stores a total of 68,000 senses for over 41,500 headwords. The majority of the senses are described by a gloss, in some cases extended by a secondary gloss. [^4] We only made use of the secondary gloss if no standard gloss was given. see Appendix B for an example of a secondary gloss. All senses are followed by example usages illustrating how the headword is used.
The two dictionaries [^5] We use the term “dictionary” for WordNet here and in the remainder of this paper after having transformed it into a dictionary-like structure (see above). represent the senses of words, describe them using both gloss and example usages, and group them by their headword. Both dictionaries differ in that their entries have gaps in different parts of their sense entries. While WordNet senses are fully covered by gloss but only have partial example usages, it is the other way around in the Swedish dictionary. They further differ in the way they treat historical senses: While WordNet only records recent senses, the Swedish dictionary covers entire modern Swedish period from 1521 onwards. Hence, the latter can be expected to cover historical senses, which fell out of use, to a much larger degree. A direct comparison of the statistics of both dictionaries is shown in Table 3.
@@ 4 Models
All our models have the same basic structure inspired by Scarlini et al. (2020), but differ in the type of data used and the embedding comparison methodology. Essentially, they first use the information given in the dictionary to create a vector representation for each existing sense. The type of information used (e.g., gloss, example usage) and how it is processed (e.g., inserting headwords if missing) are treated as model hyperparameters, as listed below. The models then create a usage embedding for the word usage to be examined and calculate the similarity to all sense embeddings of the associated headword. If the calculated similarity score between the usage embedding and the closest sense embedding is below a specific threshold, the model predicts that the word usage is the occurrence of an unknown sense and therefore marks it as unassigned. [^6] Find our code at: https://github.com/ChangeIsKey/non-recorded-sense-detection. We decided to use XL-LEXEME (see Section 2) to vectorize both target usages and senses because of its good performance in Lexical Semantic Change Detection. This task involves the detection of word sense changes and thus XL-LEXEME can be expected to encode word sense information.
Each model’s identifier is made up of a different choice of the hyperparameters, which will be explained in more detail in the following paragraphs: [^7] The total computational cost of this hyperparameter search is roughly 100 hours of GPU time.
@@ 5 Annotation
The annotators are presented instances consisting of a word usage where the target word is marked and a sense gloss that may or may not describe this usage. Additionally, they are shown the list of all possible sense glosses for the target word. The annotation approach takes inspiration from Erk et al. (2013)’s WSsim method by conducting an individual assessment of all senses for a usage, but differs in that we ask annotators for a binary classification into sense gloss fits (label “1”) and sense gloss does not fit (label “0”), instead of a five-level rating. The annotators also always have the option to choose that no specification is possible (label “-”). They are encouraged to leave a comment if this is the case. An example of an annotation instance for a usage of the word relative can be found in Appendix C.
@@ 5.1 Phase I: Random sample
In each corpus, we search word usages of a random sample of 3,000 headwords until we find at least one usage for 150 of them. We keep at most five usages for each headword, chosen at random. Lastly, we combine the word usages from the modern and the historical samples. This results in a set of word usages that is approximately equally distributed between modern and historical corpora, which can be linked to a headword in the dictionary. Combining each word usage with each eligible sense returns approximately 1,200 annotation instances, distributed over 500–700 usages, as listed in Table 5. These are assessed by the three annotators. For WordNet, only primary synsets are considered in the first annotation phase (see also Section 8).
@@
We report inter-annotator agreement scores measured by Krippendorf’s alpha (Krippendorff, 2018) in Table 7 in Appendix D. Agreement is low to moderate: In English it ranges from 0.273 to 0.561, while Swedish has more agreement (minimum 0.317, maximum 0.655). The pairwise comparison reveals that annotator A1 of English stands out in particular and the observation holds for both modern and historical usages. [^11] This annotator was excluded from the second annotation phase. For Swedish, the annotators are more in agreement, especially on the modern data.
@@ 5.2 Phase II: Model prediction
Find a summary of the annotation results in Table 6. In comparison to the first phase (see column “Unassigned baseline”), the second phase yields considerably higher shares of unassigned usages for both languages, which will be discussed in more detail in Section 6.2.
For English, the annotator-agreement on the full data is 0.384, as displayed in Table 8 in Appendix D. The agreement is comparable to the first phase. However, note that we increased the number of senses the annotators can choose from by including the number of non-primary synsets. There is considerably higher consensus among the Swedish annotators (0.56). For English, agreement on historical instances is slightly higher than on the modern instances. This contrasts with phase I. For Swedish, it is considerably worse on historical instances, as in phase I.
@@ 6.1 Phase I: Model selection
We select hyper-parameters for our models on the randomly sampled and annotated data from Phase I (see Section 5.1). In order not to overfit them on the limited data, we assess them using k-fold cross-validation. We perform a total of 10 rounds of 5-fold cross-validation for every model. Each round consists of different manipulations of the gold sense assignments. In each round, we randomly mask a set of assigned senses as unassigned in order to simulate unknown senses (cf. the method described in Erk, 2006) because the number the number of naturally unassigned senses was very low. [^12] Unassigned senses are left untouched and are included in the evaluation data. Before masking, we remove incomplete (i.e., not providing sufficient information) senses entries. Senses that are the only complete sense of a headword are always unmasked as masking them would exclude the whole target word from evaluation. We then randomly select one sense for each headword to be unmasked and mask the rest. [^13] We use the same simulated data for all models of a language, even though the gloss models in English and the example models in Swedish do not suffer from limited dictionary completeness. Find an example of this procedure described in Appendix E.
@@
Applying the randomization ten times leaves us with ten different data sets of 5-fold subdivided usages. In each round for each fold $k$, the model uses the data of the remaining four training-folds to determine a threshold $\in\{0.0,0.01,0.02,\dots,1.0\}$ (see Section 4) for the respective similarity measure, so that the $F_{0.3}$-score is maximized, giving more importance to precision over recall. [^14] Find a discussion of the evaluation metric selection in Appendix F. Using this threshold, the model then predicts on the test-fold $k$. Find a detailed example of our evaluation pipeline in Appendix E.